Consider the following differential equation:

4y(4) + y" - 18y' + 13y = et

a) Knowing that r1 = 1 is a double root, find the other two roots.

b) Find the corresponding complementary solution yc(t).

c) Find the corresponding particular solution yp(t).
Answer #1

4 M + 4 - 18y 13y=et @ auxilary equation of homogenous equation of ① uma + m² 18m +13 = 0 (9) m = 1 is a root find others by

4m² 8m +13 mail Ym3 +4m² + 5m= 13 4 m² - 4m² 2 8 m² +5m-13 8 m²-8m 13m - 13 13m-13 + el - wil Elt. ustun consider 4m² +8m +13

Roots are m=11_-1+11, -2 (b) Yo = (( + Cax) ex + e* C3 Sin (3x) + Cu Cos (3x)) complementary Solution Uc function is So UC se

x Let yp = A x² e Yp = A lx x² et 2x e y p = x² et + 2x.ex + 2xiex + 2oces + ze A ne² e AC x² 2 4x ett 2ex) + yp= c² et + a